[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Tue Oct 18, 2016

TAFB Shift Log nhc.tsbadmin at noaa.gov
Tue Oct 18 02:59:22 UTC 2016


TAFB Shift Log for Mon Oct 17 2016 22:59:19 GMT-0400 (EDT)

Shift: Mid-Shift

Pacific Forecaster: M Formosa

Atlantic Forecaster: J Lewitsky

Surface Analyst: N Ramos

Technical Issues Summary:

GFE/NCP issues:

N-AWIPS issues:

Other issues: Several Light Fixtures continue to be not working over the  
TAFB backup desk area. This problem has been ongoing for a few months. TAFB  
was previously informed that the work would have to wait until FY17.

Miscellaneous issues: Several Light Fixtures continue to be not working  
over the TAFB backup desk area. This problem has been ongoing for a few  
months. TAFB was previously informed that the work would have to wait until  
FY17.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161018/54d62a79/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list